help please! bitsat question:

help please!
bitsat question:

Grade:12

1 Answers

Yash Jain
9 years ago
The answer is 8484
See, the pattern is that, the final answer in just the repetition of a 2-digit number, so all we got to find is a 2-digit number.
Its just the simple addition. First we add up the numbers present at the bottom left of the square and we add the remaining two. If the first sum gets more than 9 than 1 is carried forward to the second sum. 
For the first square, the first sum is 7+2=9 and the other is 1+5=6. So we get the final answer as 6969. For the second square, the first sum is 7+5=12 so 2 remains and 1 is carried forward for the second sum. So the sencond sum is 4+3+1=8. Thus, the number we get is 8282.
SImilarly, third square, first sum 6+8=14 so 4 remains and 1 is carried forward. And second sum, 4+3+1=8. Hence the final answer is 8484.